how would i do this ???? 4(x-2)=x-14
Answer by Here2Help(19)

Here you go:
4(x-2)=x-14
4x-8=x-14
4x-x-8=x-14-x
3x-8=-14
3x-8+8=-14+8
3x/3=-6/3
Final answer: x=-2
Hope this helps.